WARNING: This web is not used anymore. Please use PDBService.MinuteS12July07 instead!

Minutes 28 June 07

  • Phone:
    • FNAL: Eric
    • NDGF: Olli
    • PIC: Luis
    • GRIDKA: Andreas and Doris
    • LHCb: Marco
    • TRIUMF: Denice
    • BNL: Carlos
    • RAL: Gordon

  • CERN:
    • 3D: Dirk, Eva
    • ATLAS: Florbela

  • General announcements:
    • LFC-LHCb setup completed: downstream + 4 tier1 sites (RAL, GridKA, IN2P3 and PIC). SARA has not been included yet because of last instability problems but it will be joined soon.
    • LFC team will contact Tier1 sites responsibles in order to coordinate the LFC service. To be planned in a different meeting.
    • Summary of the ATLAS activity (ATLAS week in Glasgow). Dirk will check if the presentation can be sent to the general list.
    • 2 LCG milestones for 3d fully achieved?? Florbela will check.

  • Sites status:
    • Tier 0: Setup stable. Maintenance operations on the Logminer tables.
    • SARA: new share storage has been added. Alexander is out on Wednesdays but there is one list to cover the whole week during working hours. He will send it to Dirk.
    • Triumf: patches/upgrades
    • RAL: got new storage which can be added if needed
    • BNL: got disk storage which will be added end July
    • GridKA: nothing to report
    • NDGF: Olli is going on holidays next Monday for 2 weeks. Nobody will look at the database, in case of problems, database must be removed from the Streams setup.
    • PIC: nothing to report

  • IMPORTANT: Interventions (scheduled and unscheduled) must be announced. Please use the EGEE broadcast tool and the CIC portal to announce the interventions

  • Experiments status:
    • ATLAS:
      • Scalability tests to CNAF finished: 2200 jobs
      • Numbers not definitive, wait for the COOL release which fixes some query problems, some tests will be repeated.
      • During ATLAS week: Deadline for hw request end July.
      • Conclusions about the hw requirements expected before CHEP.
      • Production schemas moved to production. Real data been replicated.
    • LHCb: nothing to report
    • CMS:
      • New release of CMS sw.
      • Some problems observed with the queries: working on it

  • Upcoming meetings.....
    • Doris has sent an abstract to Germany user group, Gordon and Dirk to the UK user group
    • Contribution to the CERN courier (Dirk and Gordon).
    • CHEP: Database session at CHEP: more focused on the experiments
    • Next workshop to plan after CHEP: PIC, new topics: procedures to add disks, ...
    • Oracle Open world: join presentation with Oracle on the 3d Streams setup.
Edit | Attach | Watch | Print version | History: r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r1 - 2007-07-17 - EvaDafonte
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    PSSGroup All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback